OpenStudy (anonymous):

Write an equation of a line that passes through the points (2,3) and (2,-6).

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]when given points (x1, y1) and (x2, y2) so plug in x1=2 , y1=3 , x2=2 , y2=-6

jigglypuff314 (jigglypuff314):

if you get a zero in the denominator, then that means the line is vertical therefore is an x = (something) then you can figure that (something) out by looking at one of the given point's x-coordinate like (2, 3) ^ x = 2